# Computational Complexity of Physical Information Sufficiency

### v3

@@ 8.3 Corollaries for Practice
[C:C8.6;R:H=succinct-hard] In the mechanized Boolean-coordinate model, existence of a sufficient set of size at most $k$ is equivalent to the relevance set having cardinality at most $k$. (L: S2P2)
@@
[C:C8.7;R:H=tractable-structured] When the bounded-action or explicit-state conditions of Theorem 6.1 hold, minimal modeling can be solved in polynomial time in the stated input size. (L: L15)
@@
This is the bounded-regime branch of Theorem 6.1, mechanized as L3.  $\blacksquare$
@@
[C:C8.8;R:H=tractable-structured] When utility is separable with explicit factors, sufficiency checking is polynomial in the explicit-state regime. (L: L16)
@@
This is the separable-utility branch of Theorem 6.1, mechanized as L4.  $\blacksquare$
@@
[C:C8.9;R:H=tractable-structured] When utility factors form a tree structure with explicit local factors, sufficiency checking is polynomial in the explicit-state regime. (L: L17)
@@
This is the tree-factor branch of Theorem 6.1, mechanized as L5.  $\blacksquare$
@@
There is a machine-checked family of reduction instances where, for non-tautological source formulas, every coordinate is relevant ($k^{*}=n$), exhibiting worst-case boundary complexity. (L: L1)
@@
The strengthened reduction proves that non-tautological source formulas induce instances where every coordinate is relevant; this is mechanized as L1.  $\blacksquare$
@@ 8.4 Hardness Distribution: Right Place vs Wrong Place
[C:P8.12;R:H=cost-growth] For any problem family $P$ measured by $H(P;n)$ above, any solution architecture $S$ and any number of use sites $n\geq 1$, if $H_{\text{total}}(S)$ is measured in the same worst-case step units over the same input family, then:
@@
For SUFFICIENCY-CHECK, Theorem 5.1 provides the baseline on the hard succinct family: $H(\text{SUFFICIENCY-CHECK};n)=2^{\Omega(n)}$ under ETH. (L: HD20)
@@
[C:P8.14;R:H=cost-growth] For fixed $n$ and positive total hardness, larger $\eta(S,n)$ is equivalent to a larger central share of realized hardness. (L: HD10)
@@
[C:C8.17;R:H=cost-growth] In the linear model, a right-hardness architecture strictly dominates a wrong-hardness architecture once use-site count exceeds central one-time hardness. Formally, for architectures $S_{\mathrm{right}},S_{\mathrm{wrong}}$ over the same problem family, if $S_{\mathrm{right}}$ has right hardness, $S_{\mathrm{wrong}}$ has wrong hardness, and $n>H_{\mathrm{central}}(S_{\mathrm{right}})$, then
@@
This is the mechanized theorem HD10.  $\blacksquare$
@@
[C:P8.18;R:H=cost-growth] This section uses two linear-model dominance modes plus generalized nonlinear dominance and boundary modes:
@@
[C:C8.19;R:H=cost-growth] For the formal native-vs-manual architecture instance, native support has lower total realized cost for all
@@
where $H_{\mathrm{baseline}}(P)$ corresponds to the Lean identifier intrinsicDOF(P) in module HardnessDistribution. (L: HD14)
@@
Immediate from HD9.  $\blacksquare$
@@ 8.5 Extension: Non-Additive Site-Cost Models
(L: HD8)
Proof.
This is exactly the mechanized theorem HD6.  $\blacksquare$
Corollary 8.23 (Eventual Generalized Dominance).
[C:C8.23;R:H=cost-growth] If condition (1) above holds and there exists $N$ such that condition (2) holds for all $m\geq N$, then there exists $N_{0}$ such that for all $n\geq N_{0}$,
@@
Immediate from HD7.  $\blacksquare$
@@
[C:P8.24;R:H=cost-growth] In the generalized model, strict right-vs-wrong dominance is not unconditional. There are explicit counterexamples:
@@
[C:T8.25;R:H=cost-growth] In the linear model of this section,
@@
[C:T8.26;R:H=cost-growth] There exists a generalized site-cost model with eventual saturation. In particular, for
@@
both $C_{K}$ and $n\mapsto H_{\text{central}}+C_{K}(n)$ are eventually saturating. (L: HD17, HD5)
@@
[C:C8.27;R:H=cost-growth] No positive-slope linear per-site model can represent the saturating family above for all $n$. (L: HD15)
@@
9 Conservation Law for Decision-Relevant Coordinates
The load-bearing fact in this section is not the set identity itself; it is the difficulty of shrinking the required set $R(P)$ in the first place. By Theorem 4.6 (and Theorem 4.7 for minimization), exact relevance identification is intractable in the worst case under succinct encoding. The identities below therefore quantify how unresolved relevance is redistributed between central and per-site work.[C:T4.6,4.7;R:S] For encoded physical systems, the same redistribution laws apply under declared physical-to-core transport: unresolved coordinates remain per-site external work, while resolved coordinates are paid once in centralized structure.[C:P2.34;T9.3,9.5,9.6;R:AR,LC] (L: CT6,HD4,HD21,HD3.)
@@ 9 Conservation Law for Decision-Relevant Coordinates
$|\text{Gap}(M,P)|+|R(P)\cap A(M)|=|R(P)|$. The total cannot be reduced, only redistributed between “handled natively” and “handled externally.” (L: HD4)
@@
This is by definition of per-site externalization and is mechanized as HD12.  $\blacksquare$
@@
such that for $n>n^{*}$, the complete model has lower total cost. (L: HD3)
@@
[C:C9.7;R:H=cost-growth] If $\text{Gap}(M,P)\neq\emptyset$, then external handling cost scales linearly with the number of decision sites. (L: HD21, HD18)
@@
The exact linear form is HD12. When the gap is nonempty (positive tax), monotone growth with $n$ is HD11.  $\blacksquare$
@@
[C:C9.8;R:H=succinct-hard] For mechanized Boolean-coordinate instances, “there exists a sufficient set of size at most kk” is equivalent to “the relevant-coordinate set has cardinality at most kk.” (L: S2P2)
@@ 10.2 Computational Decision Theory
Large-sample Bayesian-network structure learning and causal identification are already known to be hard in adjacent formulations [13, 61, 30]. Our object differs: we classify coordinate sufficiency for optimal-action invariance, with mechanized reductions and regime typing.
@@ 10.5 Energy Landscapes and Chemical Reaction Networks
Section 2.7 makes this transfer rule theorem-level: encoded physical claims are derived from core claims through an explicit assumption-transfer map, and an encoded physical counterexample refutes the corresponding core rule on that encoded slice.[C:P2.34;C2.35;R:AR](L: CT6,CT5,CT8,CT7.)
@@
Under that mapping, the explicit-vs-succinct split is the mechanized statement of access asymmetry: [E] admits polynomial verification on explicit state access, while [S+ETH] admits worst-case exponential lower bounds on succinct inputs (Theorem 5.1); finite query access already exhibits indistinguishability barriers (Proposition 5.3). The complexity separation is mechanized.[C:T5.1;P5.3;R:E,S+ETH,Q_fin](L: CC16,CC23,CC24,CC50.)
@@
The geometric core is mechanized directly: product-space slice cardinality, hypercube node-count identity, and node-vs-edge separation are proved in Lean at the decision abstraction level, with edge witnesses tied exactly to non-sufficiency.(L: DQ3, L25, L68, DP21.)
@@
The circuit-to-decision instantiation is also mechanized in a separate bridge layer: geometry and dynamics are represented as typed components of a circuit object; decision circuits add an explicit interpretation layer; molecule-level constructors instantiate both views.(L: DQ519, DQ518, DQ517, DQ155, DQ154, DQ159, DQ160.)
@@
Chemical reaction networks provide a concrete physical class where this encoding applies directly. Prior work establishes hardness for output optimization, reconfiguration, and reconstruction in CRN settings [4, 2, 19]. Those domain-specific CRN complexity results are external literature claims, not mechanized in this Lean artifact. The mechanized statement here is the transfer principle: once a CRN decision task is encoded as a C1–C4 decision problem, it inherits the same sufficiency hardness/tractability regime map (Theorems 7.2, 4.6, 6.1).[C:T7.2,4.6,6.1;R:E,S,TA](L: CR1,CC61,CC69,CC70,CC71,CC72.)
@@
The physical bridge is not a single-domain add-on: Theorem 2.37 composes law-objective semantics, reduction equivalence, and hardness lower-bound accounting in one mechanized theorem, so physical instantiation and complexity claims share one proof chain.[C:T2.37;R:AR](L: CT4.)
@@
For matter-only dynamics, the law-induced objective specialization is mechanized for arbitrary universes (arbitrary state/action types and transition-feasibility relations): the utility schema is induced from transition laws, and under a strict allowed-vs-blocked gap with nonempty feasible set, the optimizer equals the feasible-action set; with logical determinism at the action layer, the optimizer is singleton.[C:P2.36;R:AR](L: DQ8, DQ5, DQ7, DQ6.)
@@
The interface-time semantics are also explicit and mechanized: timed states are $\mathbb{N}$-indexed, one decision event is exactly one tick, run length equals elapsed interface time, and this unit-time law is invariant across substrate tags under interface-consistency assumptions.[C:P2.28,2.29,2.30,2.31;R:AR](L: DQ548,DQ537,DQ539,DQ534,DQ541,DQ549.)
@@
Regime-side interface machinery is also formalized as typed access objects with an explicit simulation preorder and certificate-upgrade statements. The access-channel and five-way composition theorems are encoded as assumption-explicit composition laws rather than informal metaphors.[C:P2.34;T2.37;R:AR](L: PC10, PC12, PC15, PC17, PC19, PC20, PC21.)
@@ 10.6 Thermodynamic Costs and Landauer’s Principle
The thermodynamic lift in Section 5 converts declared bit-operation lower bounds into declared energy/carbon lower bounds via linear conversion constants. In this artifact, those conversion consequences are mechanized conditionally through Propositions 5.28, 5.29, 5.30, and 5.31. Landauer’s principle and reversible-computation context [33, 7] define the physical assumption set; the constants are declared premises, not Lean-derived constants.[C:P5.28,5.29,5.30,5.31;R:AR](L: CC65,CC66,CC67,CC68,CC64.)
Coupling this to integrity yields a mechanized policy condition: when exact competence is hardness-blocked in the declared regime, integrity requires abstention or explicitly weakened guarantees (Propositions 3.61, 3.60). A stronger behavioral claim such as “abstention is globally utility-optimal for every objective” is not currently mechanized and remains objective-dependent.[C:P3.61,3.60;R:AR](L: IC26,IC24,CC30,IC27,IC7,IC6.)

### v4

@@ 12.3 Corollaries for Practice
[C:C12.6;R:H=succinct-hard] In the mechanized Boolean-coordinate model, existence of a sufficient set of size at most $k$ is equivalent to the relevance set having cardinality at most $k$. (L: S2P2)
@@
[C:C12.7;R:H=tractable-structured] When the bounded-action or explicit-state conditions of Theorem 10.1 hold, minimal modeling can be solved in polynomial time in the stated input size. (L: CC81)
@@
This is the bounded-regime branch of Theorem 10.1, mechanized as CC81.  $\blacksquare$
@@
[C:C12.8;R:H=tractable-structured] When utility is separable with explicit factors, sufficiency checking is polynomial in the explicit-state regime. (L: CC82)
@@
This is the separable-utility branch of Theorem 10.1, mechanized as CC82.  $\blacksquare$
@@
[C:C12.9;R:H=tractable-structured] When utility factors form a tree structure with explicit local factors, sufficiency checking is polynomial in the explicit-state regime. (L: CC84)
@@
This is the tree-factor branch of Theorem 10.1, mechanized as CC84.  $\blacksquare$
Corollary 12.10 (Low-Tensor-Rank Tractability).
[C:C12.10;R:H=tractable-structured] When utility has tensor rank $R$ with explicit factor decomposition $U(a,s)=\sum_{r=1}^{R}w_{r}f_{r}(a)\prod_{i=1}^{n}g_{ri}(s_{i})$, sufficiency checking is polynomial in $|A|\cdot R\cdot n$. (L: CC71)
Proof.
This is the low-tensor-rank branch of Theorem 10.1, mechanized as CC71.  $\blacksquare$
Corollary 12.11 (Bounded-Treewidth Tractability).
[C:C12.11;R:H=tractable-structured] When utility decomposes as pairwise interactions on a graph of treewidth $w$, sufficiency checking is polynomial in $n\cdot k^{w+1}$. (L: CC73, CC75)
Proof.
This is the bounded-treewidth branch of Theorem 10.1, mechanized as CC73.  $\blacksquare$
Corollary 12.12 (Coordinate-Symmetry Tractability).
[C:C12.12;R:H=tractable-structured] When state space is $S=[k]^{d}$ with coordinate-permutation-invariant utility, sufficiency checking is polynomial in $\binom{d+k-1}{k-1}^{2}\cdot|A|$, which is polynomial in $d$ for fixed $k$. (L: CC72)
Proof.
This is the coordinate-symmetry branch of Theorem 10.1, mechanized as CC72.  $\blacksquare$
@@
There is a machine-checked family of reduction instances where, for non-tautological source formulas, every coordinate is relevant ($k^{*}=n$), exhibiting worst-case boundary complexity. (L: CC26)
@@
The strengthened reduction proves that non-tautological source formulas induce instances where every coordinate is relevant; this is mechanized as CC26.  $\blacksquare$
@@ 12.4 Hardness Distribution: Right Place vs Wrong Place
[C:P12.15;R:H=cost-growth] For any problem family $P$ measured by $H(P;n)$ above, any solution architecture $S$ and any number of use sites $n\geq 1$, if $H_{\text{total}}(S)$ is measured in the same worst-case step units over the same input family, then:
@@
For SUFFICIENCY-CHECK, Theorem 9.1 provides the baseline on the hard succinct family: $H(\text{SUFFICIENCY-CHECK};n)=2^{\Omega(n)}$ under ETH. (L: HD23, HD25)
@@
[C:P12.17;R:H=cost-growth] For fixed $n$ and positive total hardness, larger $\eta(S,n)$ is equivalent to a larger central share of realized hardness. (L: HD11)
@@
[C:C12.20;R:H=cost-growth] In the linear model, a right-hardness architecture strictly dominates a wrong-hardness architecture once use-site count exceeds central one-time hardness. Formally, for architectures $S_{\mathrm{right}},S_{\mathrm{wrong}}$ over the same problem family, if $S_{\mathrm{right}}$ has right hardness, $S_{\mathrm{wrong}}$ has wrong hardness, and $n>H_{\mathrm{central}}(S_{\mathrm{right}})$, then
@@
This is the mechanized theorem HD19.  $\blacksquare$
@@
[C:P12.21;R:H=cost-growth] This section uses two linear-model dominance modes plus generalized nonlinear dominance and boundary modes:
@@
[C:C12.22;R:H=cost-growth] For the formal native-vs-manual architecture instance, native support has lower total realized cost for all
@@
where $H_{\mathrm{baseline}}(P)$ corresponds to HD15. (L: HD15)
@@
Immediate from HD15.  $\blacksquare$
@@ 12.5 Extension: Non-Additive Site-Cost Models
This is exactly the mechanized theorem HD9.  $\blacksquare$
Corollary 12.26 (Eventual Generalized Dominance).
[C:C12.26;R:H=cost-growth] If condition (1) above holds and there exists $N$ such that condition (2) holds for all $m\geq N$, then there exists $N_{0}$ such that for all $n\geq N_{0}$,
(L: HD10)
Proof.
Immediate from HD10.  $\blacksquare$
@@
[C:P12.27;R:H=cost-growth] In the generalized model, strict right-vs-wrong dominance is not unconditional. There are explicit counterexamples:
@@
[C:T12.28;R:H=cost-growth] In the linear model of this section,
@@
[C:T12.29;R:H=cost-growth] There exists a generalized site-cost model with eventual saturation. In particular, for
@@
both $C_{K}$ and $n\mapsto H_{\text{central}}+C_{K}(n)$ are eventually saturating. (L: HD20, HD6)
@@
[C:C12.30;R:H=cost-growth] No positive-slope linear per-site model can represent the saturating family above for all $n$. (L: HD16)
@@
13 Redistribution Identity for Decision-Relevant Coordinates
Formally: this section defines expressive-gap tax and proves redistribution, growth, and amortization identities under fixed required-coordinate sets.
The load-bearing fact in this section is not the set identity itself; it is the difficulty of shrinking the required set $R(P)$ in the first place. By Theorem 4.6 (and Theorem 4.7 for minimization), exact relevance identification is intractable in the worst case under succinct encoding. The identities below therefore quantify how unresolved relevance is redistributed between central and per-site work.[C:T4.6, 4.7;R:S] For encoded physical systems, the same redistribution laws apply under physical-to-core transport: unresolved coordinates remain per-site external work, while resolved coordinates are paid once in centralized structure.[C:P2.42;T13.3, 13.5, 13.6;R:AR, LC] (L: CT6, HD5, HD26, HD21, HD4.)
@@ 13 Redistribution Identity for Decision-Relevant Coordinates
$|\text{Gap}(M,P)|+|R(P)\cap A(M)|=|R(P)|$. The total cannot be reduced, only redistributed between “handled natively” and “handled externally.” (L: HD5)
@@
This is by definition of per-site externalization and is mechanized as HD26.  $\blacksquare$
@@
such that for $n>n^{*}$, the complete model has lower total cost. (L: HD4)
@@
[C:C13.7;R:H=cost-growth] If $\text{Gap}(M,P)\neq\emptyset$, then external handling cost scales linearly with the number of decision sites. (L: HD26, HD21)
@@
The exact linear form is HD26. When the gap is nonempty (positive tax), monotone growth with $n$ is HD21.  $\blacksquare$
@@
[C:C13.8;R:H=succinct-hard] For mechanized Boolean-coordinate instances, “there exists a sufficient set of size at most kk” is equivalent to “the relevant-coordinate set has cardinality at most kk.” (L: S2P2)
@@
Informally: when exact minimization is hard, unresolved relevance shifts between centralized structure and repeated local work.
@@ 14.2 Computational Decision Theory
Large-sample Bayesian-network structure learning and causal identification are already known to be hard in adjacent formulations [17, 80, 41]. Our object differs: we classify coordinate sufficiency for optimal-action invariance across static/stochastic/sequential regimes, with theorem-level placements and mechanized reductions.
@@
Functional information and selection.
Prior work defines function-relative information objects in physical systems. Szostak defines functional information as the negative log of the fraction of configurations achieving a specified function [87]. Wong et al. study function and selection in evolving systems and state an increasing-functional-information law under selection [93]. This paper addresses a complementary object: the computational decision predicate for when projected coordinates are sufficient to preserve optimal-action correspondence, with explicit complexity placements and mechanized proofs.
@@ 14.5 Energy Landscapes and Chemical Reaction Networks
Section 2.7 makes this transfer rule theorem-level: encoded physical claims are derived from core claims through an explicit assumption-transfer map, and an encoded physical counterexample refutes the corresponding core rule on that encoded slice.[C:P2.42;C2.43;R:AR](L: CT6, CT5, CT8, CT7.)
@@
Under that mapping, the explicit-vs-succinct split is the mechanized statement of access asymmetry: [E] admits polynomial verification on explicit state access, while [S+ETH] admits worst-case exponential lower bounds on succinct inputs (Theorem 9.1); finite query access already exhibits indistinguishability barriers (Proposition 9.3). The complexity separation is mechanized.[C:T9.1;P9.3;R:E, S+ETH, Q_fin](L: CC16, CC23, CC24, CC50.)
@@
The geometric core is mechanized directly: product-space slice cardinality, hypercube node-count identity, and node-vs-edge separation are proved in Lean at the decision abstraction level, with edge witnesses tied exactly to non-sufficiency.(L: DG5, DG6, DG15, DG16, DG18.)
@@
The circuit-to-decision instantiation is also mechanized in a separate bridge layer: geometry and dynamics are represented as typed components of a circuit object; decision circuits add an explicit interpretation layer; molecule-level constructors instantiate both views.(L: IN4, IN5, IN6, IN13, IN14, IN15, IN16.)
@@
Chemical reaction networks provide a concrete physical class where this encoding applies directly. Prior work establishes hardness for output optimization, reconfiguration, and reconstruction in CRN settings [4, 2, 26]. Those domain-specific CRN complexity results are external literature claims, not mechanized in this Lean artifact. The mechanized statement here is the transfer principle: once a CRN decision task is encoded as a C1–C4 decision problem, it inherits the same sufficiency hardness/tractability regime map (Theorems 11.2, 4.6, 10.1).[C:T11.2, 4.6, 10.1;R:E, S, TA](L: CR1, CC61, CC69, CC70, CC71, CC72.)
@@
The physical bridge is not a single-domain add-on: Theorem 2.45 composes law-objective semantics, reduction equivalence, and hardness lower-bound accounting in one mechanized theorem, so physical instantiation and complexity claims share one proof chain.[C:T2.45;R:AR](L: CT4.)
@@
For matter-only dynamics, the law-induced objective specialization is mechanized for arbitrary universes (arbitrary state/action types and transition-feasibility relations): the utility schema is induced from transition laws, and under a strict allowed-vs-blocked gap with nonempty feasible set, the optimizer equals the feasible-action set; with logical determinism at the action layer, the optimizer is singleton.[C:P2.44;R:AR](L: DQ8, DQ5, DQ7, DQ6.)
@@
The interface-time semantics are also explicit and mechanized: timed states are $\mathbb{N}$-indexed, one decision event is exactly one tick, run length equals elapsed interface time, and this unit-time law is invariant across substrate tags under interface-consistency assumptions.[C:P2.34, 2.35, 2.36, 2.37;R:AR](L: DT6, DT8, DT10, DT13, DT16, DT24.)
@@
Regime-side interface machinery is also formalized as typed access objects with an explicit simulation preorder and certificate-upgrade statements. The access-channel and five-way composition theorems are encoded as assumption-explicit composition laws.[C:P2.42;T2.45;R:AR](L: ARG2, ARG3, ARG6, ARG13, ARG17, ARG19, ARG12.)
@@ 14.6 Thermodynamic Costs and Landauer’s Principle
The thermodynamic lift in Section 9 converts bit-operation lower bounds into energy/carbon lower bounds via linear conversion constants. In this artifact, those conversion consequences are mechanized with explicit premises through Propositions 9.28, 9.29, 9.30, and 9.31. Landauer’s principle and reversible-computation context [44, 8] define the physical assumption set; the constants are premises, not Lean-derived constants.[C:P9.28, 9.29, 9.30, 9.31;R:AR](L: CC65, CC66, CC67, CC68, CC64.)
Physics assumptions decompose into regime tags.
The paper does not create domain-specific tags for physics assumptions. Instead, every physics assumption decomposes into the same regime tags used throughout the paper. This uniformity states that physics assumptions are not epistemically privileged; they are declared assumptions, tracked with the same discipline as the paper’s complexity-theoretic assumptions. The explicit physical-to-core mapping does not make the results less physical. It standardizes premise exposure and transfer obligations: accepted physics assumptions and this model are handled under the same declared-assumption framework, then checked with the same mechanized proof discipline where formalized.
Table 1: Physics-Assumption-to-Regime-Tag Decomposition
Physics Assumption Regime Decomposition Justification
Landauer bound ($E\geq kT\ln 2$) [AR], [H=cost-growth] Declared law; yields cost lower bounds
Lorentz invariance [AR] Declared physical symmetry
Discrete state space [Q_fin], [AR] Finite-state core; declared assumption
Thermodynamic constants ($\lambda,\rho$) [AR], [H=cost-growth] Declared conversion model
Planck scale ($\ell_{P}$) [AR] Declared physical constant
Physical-to-core encoding ($E$) [AR] Declared encoding map
The decomposition in Table 1 makes explicit what the paper already practices: physics assumptions are declared premises in the [AR] (axiomatic regime) sense, and when they yield cost lower bounds, they inherit [H=cost-growth]. The finite-state query core [Q_fin] captures the discrete-state assumption underlying the computational model. No physics assumption receives a special domain tag; all decompose into the existing regime taxonomy. This is a standardization claim about proof form, not a change in physical content.[C:P9.28, 9.52, 9.34, 9.33;T2.45;R:AR, H=cost-growth, Q_fin]
Coupling this to integrity yields a mechanized policy condition: when exact competence is hardness-blocked in the active regime, integrity requires abstention or explicitly weakened guarantees (Propositions 3.61, 3.60). A stronger behavioral claim such as “abstention is globally utility-optimal for every objective” is not currently mechanized and remains objective-dependent.[C:P3.61, 3.60;R:AR](L: IC26, IC24, CC30, IC27, IC7, IC6.)
14.7 Neukart–Vinokur Thermodynamic Duality
Neukart and Vinokur [56] propose a thermodynamic duality $dU=T\,dS-p\,dV+\lambda\,dC$ where $C$ is a “complexity coordinate” associated with computational irreversibility in spin-glass systems. Corollary 9.32 derives their duality as a specialization of our thermodynamic lift (Proposition 9.28) for $\lambda>0$. [C:C9.32;P9.28;R:AR] (L: CC65, CC50.)
Their spin-glass phase transitions become regime-specific instances: our [S+ETH] hard families with $k^{*}=n$ witness their “complexity potential” via explicit TAUTOLOGY reduction (Theorem 9.1). Physical transport via $E:\mathcal{P}\to\mathcal{D}$ extends to arbitrary substrates, not only quantum or spin systems (Theorem 2.45). [C:T9.1, 2.45;R:S+ETH, AR] (L: CC16, CT4.)
Thus Neukart–Vinokur $\subseteq$ [AR] regime: $\lambda,\rho>0$, explicit encoding $E:\mathcal{P}\to\mathcal{D}$ $\Rightarrow$ their thermodynamic duality holds by assumption transfer through our coNP-complete core (Proposition 2.42). [C:P2.42;R:AR] (L: CT6.)